How many atoms do mercury vapour molecules consist of if the density of mercury vapour relative to air is 6.92.The average mass of air is 29g/mol
1
Expert's answer
2017-03-26T04:51:06-0400
D air (mercury vapour) = M(mercury vapour) / M(air)
M(mercury vapour) = D air (mercury vapour)*M(air) = 6.92*29 = 200.68 (g/mol)
M(Hg) = 200.6 g/mol that’s why mercury vapour molecules consist of 1 atom

Answer: mercury vapour molecules consist of 1 atom